OpenStudy (anonymous):

Express the inverse of f(x)=6x-1 in the form y = f^-1(x)

OpenStudy (anonymous):

Inverse: Take the x and the y and switch them. We can assume that f(x) = y and therefore the equation can be written as: y=6x-1 Swapping the x and y gives us: x=6y-1 Now solve for y.

OpenStudy (anonymous):

y=1/6 + x/6?

jimthompson5910 (jim_thompson5910):

Similar Example: Find the inverse of \[\Large f(x) = 9x + 20\] Basic Idea: Replace f(x) with y. Swap x and y, then solve for y. After getting y isolated, replace y with \(\Large f^{-1}(x)\) to get the inverse function. \[\Large f(x) = 9x + 20\] \[\Large y = 9x + 20\] \[\Large x = 9y + 20\] \[\Large x - 20 = 9y\] \[\Large 9y = x - 20\] \[\Large y = \frac{x - 20}{9}\] \[\Large f^{-1}(x) = \frac{x - 20}{9}\] So the inverse function is \[\Large f^{-1}(x) = \frac{x - 20}{9}\] Keep in mind that this is NOT your answer as it's only an example.

OpenStudy (anonymous):

f^(-1) (x) = x+1/6?

jimthompson5910 (jim_thompson5910):

if you mean \[\Large f^{-1}(x) = \frac{x+1}{6}\] then you are 100% correct

OpenStudy (anonymous):

yes!!!

OpenStudy (anonymous):

:)